A bond has a $1,000 face value and a $1,146 market value. The bond pays interest semi-annually, has a yield-to-maturity of 7.47 percent, and matures in 10 years. What is the current yield?
Answer #1

Rate = 7.47% / 2 = 3.735%

Number of periods = 10 * 2 = 20

Price = Coupon * [1 - 1 / (1 + r)n] / r + Fv / (1 + r)n

1146 = Coupon * [1 - 1 / (1 + 0.03735)20] / 0.03735 + 1000 / (1 + 0.03735)20

1146 = Coupon * [1 - 0.480279] / 0.03735 + 480.279195

1146 = Coupon * 13.914886 + 480.279195

Coupon = 47.8424

Annual coupon = 47.8424 * 2 = 95.6848

Current yield = (Annual coupon / price) * 100

Current yield = (95.6848 / 1146) * 100

Current yield = 9.57%